OpenStudy (anonymous):

Which of the following is equivalent to six and three-sevenths divided by twelve

OpenStudy (anonymous):

@jim_thompson5910 help?

jimthompson5910 (jim_thompson5910):

six and three-sevenths = 6 & 3/7 6 & 3/7 = 6/1 + 3/7 6 & 3/7 = (6/1)*(7/7) + 3/7 6 & 3/7 = 42/7 + 3/7 6 & 3/7 = (42 + 3)/7 6 & 3/7 = 45/7 making sense so far?

OpenStudy (anonymous):

Yes but your not saying thats the answer right? because my answer choices are... 6 1/28 1 13/15 15/28 2 1/28

jimthompson5910 (jim_thompson5910):

we're not done, but is it making sense so far?

OpenStudy (anonymous):

@jim_thompson5910 thats what I thought and yes!

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

6 & 3/7 = 45/7 so saying " six and three-sevenths divided by twelve" is the same as saying "6 & 3/7 divided by 12" and that's the same as writing this "(45/7) divided by 12"

OpenStudy (anonymous):

so I divided 45/7 divided by 12?

jimthompson5910 (jim_thompson5910):

(45/7) divided by 12 = (45/7)/(12/1) (45/7) divided by 12 = (45/7)*(1/12) (45/7) divided by 12 = (45*1)/(7*12) (45/7) divided by 12 = 45/84 (45/7) divided by 12 = 15/28

jimthompson5910 (jim_thompson5910):

to reduce in that last step, you divide each piece by the GCF 3

OpenStudy (anonymous):

so I divided each answer choice by 3?

OpenStudy (anonymous):

nvm so the answer is 15/28 right?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

@jim_thompson5910 thanks so much!

jimthompson5910 (jim_thompson5910):

yw